A QUANTUM CHEMICAL INVESTIGATION OF THE STABILITY AND CHEMISTRY OF THE ANIONS OF CO AND H_{2}CO IN ASTROPHYSICAL ICES

\underline{L. CHEN} and D. E. WOON, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana IL 61801.

Electrons may either be deposited on interstellar icy grain mantles from the gas phase, or they may be produced in situ via photoionization. It is energetically favorable for electrons to bind to species in the ice to form anions. Similarly, anions may be formed within icy grain mantles from chemical processes such as proton transfer from an acid to a base. To understand the fate of electrons in astrophysical ices and the potential impact of anions on gas-grain chemistry, we explored the behavior of two common and important interstellar molecules – CO and formaldehyde (H_{2}CO) – when they interact with an electron in ice. Neither CO nor H_{2}CO bind an electron in the gas phase. In astrophysical ices, however, quantum chemical calculations indicate that the anions of CO and H_{2}CO are stabilized by interacting with the polarized field provided by water ice and are bound with respect to CO and H_{2}CO. We then investigated whether CO^{−} and H_{2}CO^{−} are stable with respect to subsequent chemical reactions and sought to identify vibrational spectroscopic features that might be useful for experimentalists to identify these anions in the laboratory. In our theoretical investigations reported here, we studied CO^{−} + nH_{2}O and H_{2}CO^{−} + mH_{2}O in clusters containing from 2 to 17 water molecules using B3LYP/6-31+G^{∗∗} and MP2/6-31+G^{∗∗} calculations. The calculations show that in smaller water clusters, both CO^{−} and H_{2}CO^{−} are stabilized by water molecules. However, in larger clusters, both anions can undergo barrierless reactions in which a proton is transferred from a nearby water molecule, forming radicals such as HCO or H_{2}COH and hydroxide anion (OH^{−}). Since HCO and H_{2}COH are precursors to H_{2}CO and CH_{3}OH, electron-induced anion chemistry in ices may contribute to the observed abundances of important astromolecules. In addition to the structures and energetics of these clusters, we will also provide theoretical predictions of spectroscopic features that might be detected in the laboratory or interstellar medium.